40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2800     Entry time: Tue Apr 13 20:02:02 2010
Author: Koji 
Type: Update 
Category: SUS 
Subject: BS chamber opened, PRM/SRM SOS removed from the table 

Bob, Steve, and Koji

We opened North heavy door of the BS chamber in the afternoon.

In the evening, Koji worked on the PRM/SRM removal.

- Cleaned up the OPLEV mirrors to create some spaces near the door.

- Clamped PRM/SRM.

- Removed OSEMs. Made a record of the OSEMs. The record is on the wiki (http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/Suspensions)

- Found the SOSs are quite easy to remove from the table as they are shorter than the MOSs.

- Put a new Al sheet on a wagon. Put the SOSs on it. Wrapped them by the Al foils.

- Carried it to the clean room. They are on the right flow bench. Confirmed the wires are still fine.

- Closed up the chamber putting a light door.

Attachment 1: IMG_2384.jpg  81 kB  | Show | Show all
ELOG V3.1.3-